OpenStudy (anonymous):

tracy needs to find the length and width of the equipment storage room where she works. The room is 2 meters than it is wide. Its perimeter is 24 meters.

OpenStudy (0_0youscareme):

Perimeter is the sum of all sides. A rectangle has 4 sides, 2 width and 2 length. So, P=24=2W+2L We also know that the length and width can be related by L=W+2 Replace L with W+2 in the original equation. P=24=2W+2L=2W+2(W+2)=4W+4 Solve for W 20=4W W=5 L=7